Calculate the magnitudes of magnetic intensity vector H and magnetic field vector B at the

centre of a 2500-turn solenoid which is 0.50 m long and carries a current of 1.0 A.

(uo= 4π x 10-7Hm-1)



1
Expert's answer
2021-06-09T08:06:01-0400

Magnetic field at the center of a solenoid

"B=\frac{\mu_0NI}{L}"

"B=\frac{4\pi.10^{-7}\frac{Tm}{A}.2500.\space1.0A}{0.5m}=6.28\space ^.10^{-3}T"

Magnitude of magnetic intensity:

"B=\mu_0H,"

"H=\frac{B}{\mu_0}=\frac{6.28\space^.10^{-3}T}{4\pi ^.10^{-7} \frac{Tm}{A}}=4997\frac{A}{m}."


"B=6.28^.10^{-3}T,H=4997 \frac{A}{m}."